Chapter 13 question: 1. A bakery buys flour in 25-pound bags. The bakery uses 1,215 bags a year. Ordering cost is $20 per order. Annual carrying cost is $150 per bag. a) Determine the economic order quantity. b) What is the average number of bags on hand? c) How many orders per year will there be? d) Compute the total cost of ordering and carrying flour.
Answer #1

Solution:

Annual demand (D) = 1,215 bags a year

Ordering cost (Co) = $20 per order

Annual carrying cost (Cc) = $150 per bag

(a) Economic Order Quantity (EOQ):

EOQ = SQRT [(2 x D x Co) / Cc]

EOQ = SQRT [(2 x 1,215 x $20) / $150]

EOQ = 18

Economic Order Quantity = 18 bags

(b) Average number of bags on hand:

Average number of bags = EOQ / 2

Average number of bags = 18 / 2

Average number of bags = 9 bags

(c) Number of Orders:

Number of orders = Annual demand / EOQ

Number of orders = 1,215 / 18

Number of orders = 67.5 orders

(d) Total cost of ordering and carrying:

Total cost = Ordering cost + Carrying cost

Total cost = [(D / EOQ) x Co] + [(EOQ / 2) x Cc]

Total cost =  [(1,215 / 18) x $20] + [(18 / 2) x $150]

Total cost = $1,350 + $1,350

Total cost = $2,700